# Quantum Dynamics and Information Measures in $\mathcal{P}$$\mathcal{T}$ and Anti - $\mathcal{P}$$\mathcal{T}$-Symmetric Systems

### v1

@@ I Introduction
These symmetry conditions are particularly relevant in open quantum systems, where non-Hermiticity naturally arises due to environmental coupling [2]. $\mathcal{P}$$\mathcal{T}$-symmetric systems display a well-known phase transition between an unbroken phase, with real eigenvalues and quasi-unitary evolution, and a broken phase, where eigenvalues appear in complex-conjugate pairs and probability is not conserved [3, 4]. In Anti - $\mathcal{P}$$\mathcal{T}$-symmetric systems, the dynamics are inherently dissipative but exhibit distinct coherence patterns that can be harnessed in engineered quantum setups [5, 6].
Recent research has unveiled a broad spectrum of applications for $\mathcal{P}$$\mathcal{T}$ and Anti - $\mathcal{P}$$\mathcal{T}$-symmetry in diverse quantum platforms such as quantum computing  [7, 8, 9, 10], quantum sensing  [11, 6, 12, 13], quantum optics and photonics  [14, 15, 16, 17], quantum thermodynamic  [18, 19, 20], quantum chaos and scrambling  [21, 22] and topological quantum systems  [23, 24].
@@ III.2 Dynamics of $\mathcal{P}$$\mathcal{T}$-symmetric qubit
First, consider a Hermitian system characterized by a diagonal Hermitian Hamiltonian $h_{D}$, with an associated density matrix $\rho_{Dh}$. The standard Liouville–Von Neumann equation for this setup is:
@@
The reduced density matrix for the system under Anti - $\mathcal{P}$$\mathcal{T}$-symmetry evolves as:
@@ Conclusion
Anti - $\mathcal{P}$$\mathcal{T}$-symmetric systems, where $\{PT,H\}=0$, behave similar the $\mathcal{P}$$\mathcal{T}$-symmetric systems. However, the key difference is that in Anti - $\mathcal{P}$$\mathcal{T}$systems, $\xi$ and $\delta$ are the non-Hermitian parameters and govern the decoherence behavior.
@@
Overall, non-Hermitian systems with $\mathcal{P}$$\mathcal{T}$and Anti - $\mathcal{P}$$\mathcal{T}$symmetry tend to be more applicable in quantum information and related fields, for example quantum cryptography or computation, due to the their behavior. Systems with Anti - $\mathcal{P}$$\mathcal{T}$symmetry outshine $\mathcal{P}$$\mathcal{T}$-symmetric systems in performance under open-system dynamics and decoherence, paving the way for more reliable resources for further applications.
@@ Acknowledgments
We would like to thank Mina Ghodsi Yengejeh and Dr. Abasalt Rostami for thier valuable support and comments to this work.

### v2

@@ I Introduction
These symmetry conditions are particularly relevant in open quantum systems, where non-Hermiticity naturally arises due to environmental coupling [2]. $\mathcal{P}$$\mathcal{T}$-symmetric systems display a well-known phase transition between an unbroken phase, with real eigenvalues and quasi-unitary evolution, and a broken phase, where eigenvalues appear in complex-conjugate pairs and probability is not conserved [3, 4]. In Anti - $\mathcal{P}$$\mathcal{T}$-symmetric systems, the dynamics are inherently dissipative but exhibit distinct coherence patterns that can be harnessed in engineered quantum setups [5, 6, 7].
Recent research has unveiled a broad spectrum of applications for $\mathcal{P}$$\mathcal{T}$ and Anti - $\mathcal{P}$$\mathcal{T}$-symmetry in diverse quantum platforms such as quantum computing  [8, 9, 10, 11], quantum sensing  [12, 6, 13, 14], quantum optics and photonics  [15, 16, 17, 18, 19], quantum thermodynamic  [20, 21, 22], quantum chaos and scrambling  [23, 24] and topological quantum systems  [25, 26].
@@ III.2 Dynamics of $\mathcal{P}$$\mathcal{T}$-symmetric qubit
First, consider a Hermitian system characterized by a diagonal Hermitian Hamiltonian $h_{D}$, with an associated density matrix $\rho_{Dh}$. Putting $\hbar$=1, the standard Liouville–Von Neumann equation for this setup is:
@@
[…]
@@ Conclusion
Anti - $\mathcal{P}$$\mathcal{T}$-symmetric systems, where $\{PT,H\}=0$, behave similar the $\mathcal{P}$$\mathcal{T}$-symmetric systems. However, the key difference is that in Anti - $\mathcal{P}$$\mathcal{T}$systems, $\xi$ and $\delta$ are the non-Hermitian parameters and govern the decoherence behavior [5].
@@
Overall, non-Hermitian systems with $\mathcal{P}$$\mathcal{T}$and Anti - $\mathcal{P}$$\mathcal{T}$symmetry tend to be more applicable in quantum information and related fields, for example quantum cryptography or computation, due to the their behavior. Systems with Anti - $\mathcal{P}$$\mathcal{T}$symmetry outshine $\mathcal{P}$$\mathcal{T}$-symmetric systems in performance under open-system dynamics and decoherence [5, 7], paving the way for more reliable resources for further applications.
@@ Acknowledgments
We would like to thank Mina Ghodsi Yengejeh and Abasalt Rostami for thier valuable support and comments to this work.